Instead of measuring resistance suppose that in (a) data of the variation of
current with the length of wire is collected. Show how such data can be
used to determine the resistivity of the wire
1
Expert's answer
2019-03-01T10:58:47-0500

Applying Ohm's law derive an expression for current:


"I=\frac{V}{R}=\frac{V}{\rho l/A},"

and express the resistivity:


"\rho=\frac{VA}{Il}."

Since in experiments most probably you know the voltage "V" and the cross-sectional area "A" of the conductor, just substitute corresponding "I" and "l" from the point on the graph.